Please help me write the slope - intercept equation for the line with the given slope that contains the given point. m=-3;(-2,1)
now I know my formula will be y=mx+b

Answer by mananth(16949)

m=-3

Plug value of the slope and point (-2 ,1) in
Y=mx+b
1=6+b
b=1-6
b=-5
So the equation will be
Y =-3x-5